OpenStudy (anonymous):

Find the equation of the circle in center radius from given endpoints of a diameter (3,12) and (9,4)

OpenStudy (jdoe0001):

ttp://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/general-formula-equation-of-circle.png

OpenStudy (jdoe0001):

http://www.mathwarehouse.com/geometry/circle/images/equation-of-circle/general-formula-equation-of-circle.png those 2 points are the diameter.... so the center is half-way between them so \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(3\quad ,&12)\quad &(9\quad ,&4) \end{array} \\\quad \\ \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\Longleftarrow \color{red}{center, (h, k)}\\ \\ \quad \\ \text{distance between 2 points}\\ d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\quad \Longleftarrow \color{red}{diameter}\\ \quad \\ radius = \cfrac{diameter}{2} \)